OpenStudy (anonymous):

Help? Find all solutions in the interval [0, 2π). sin2x - cos2x = 0

OpenStudy (anonymous):

sin2x = cos2x sin2x = sin(pi/2-2x) is last step clear to you?

OpenStudy (anonymous):

Oh, crud. I'm sorry. The equation is: \[\sin^2 x = \cos^2 x\]

OpenStudy (phi):

you could do a few things: use sin^2 = 1 - cos^2 and put everything in terms of cos (or sin) or divide by cos^2 and solve \(\tan^2 x =1\)

OpenStudy (anonymous):

@This_Is_Batman \[\sin^2x-\cos^2x = 0\] \[\cos^2x-\sin^2x=0\] we know that \[\cos^2x-\sin^2x = \cos2x\] therefore, \[\cos2x = 0 \] is all steps clear to you?

OpenStudy (anonymous):

Uhmm... I think so..

OpenStudy (anonymous):

How do I find the solutions????

OpenStudy (anonymous):

\[\cos2x = \cos(\frac{ n \pi }{ 2 })\] where n is odd number \[2x = \frac{ n \pi }{ 2}\] \[x=\frac{ n \pi }{4 }\]

OpenStudy (anonymous):

What?!

OpenStudy (anonymous):

do u know how to make cosine graph?

OpenStudy (anonymous):

 graph is not drawn very continuously apologies for that from the graph it can be concluded that cosine is zero at pi/2 , 3pi/2, 5pi/2 and so on so i write it down as a general formula cos(npi/2) where n= 1,3,5,7,so on